OpenStudy (anonymous):

64^(2-x)=4^(4x)... can someone please help me and explain how to solve this? thank you :)

OpenStudy (asnaseer):

use the fact:\[64=4^3\]

OpenStudy (amistre64):

the rest is just loggy

OpenStudy (asnaseer):

you don't need to use logs

OpenStudy (amistre64):

you do if you need the practice lol

OpenStudy (asnaseer):

you should end up with 4^{something} = 4^{something-else} then just equate the powers

OpenStudy (amistre64):

equate exponents is the same as logging

OpenStudy (anonymous):

when u smarties are done here, come my way! haha

OpenStudy (asnaseer):

sunshine - do you understand how to solve this now?

OpenStudy (anonymous):

no wait I am confused :/

OpenStudy (anonymous):

i am trying to find the solution but idk how.

OpenStudy (amistre64):

divide each side by 4^3 might plausible

OpenStudy (asnaseer):

you are given this:\[64^{2-x}=4^{4x}\]now replace the 64 with \(4^3\) to get:\[(4^3)^{2-x}=4^{4x}\]therefore:\[4^{3(2-x)}=4^{4x}\]

OpenStudy (amistre64):

yep, then ignore the 4s and equate the exponential bits

OpenStudy (anonymous):

Oh!And then what would I have to do?

OpenStudy (anonymous):

Ignore all 4s?

OpenStudy (asnaseer):

if you are given:\[4^a=4^b\]what does that tell you about a and b?

OpenStudy (amistre64):

lol, it makes no sense to ignore ALL the 4s, just the useless ones ;)

OpenStudy (asnaseer):

sunshine: can you answer my previous question above?

OpenStudy (anonymous):

Does it mean they're the same? I am sorry I am not positive

OpenStudy (amistre64):

\[log_4[4^a=4^b]\] \[log_4(4^a)=log_4(4^b)\] \[a\ log_4(4)=b\ log_4(4)\] \[\frac{a\ \cancel{log_4(4)}}{\cancel{log_4(4)}}=\frac{b\ \cancel{log_4(4)}}{\cancel{log_4(4)}}\] \[a=b\]

OpenStudy (asnaseer):

yes - thats right - it means a=b so use this fact to work out the solution to your problem.

OpenStudy (asnaseer):

remember we ended up with:\[4^{3(2-x)}=4^{4x}\]

OpenStudy (anonymous):

Alright, Thank you!!!

OpenStudy (asnaseer):

yw

OpenStudy (anonymous):

IS the answer 6/7?

OpenStudy (asnaseer):

yup - well done!